# Elementary Quantum Gates from Lie Group Embeddings in $U(2^{n})$: Geometry, Universality, and Discretization

### v1

@@ Abstract
In the standard circuit model, elementary gates are specified relative to a chosen tensor factorization and are therefore extrinsic to the ambient group $U(2^{n})$. Writing $N=2^{n}$, we introduce an intrinsic descriptor layer in $U(N)$ by declaring as primitive the motions inside faithful embedded copies of $SU(2)$, leading to the phase-free dictionary $\mathcal{G}^{SU}_{\mathrm{elem}}(n)=\bigcup_{\phi\in\mathrm{Emb}(SU(2),U(N))}\phi(SU(2))$, and we also discuss the phase-inclusive $U(2)$ variant. We show that $\mathrm{Emb}(SU(2),U(N))$ decomposes into finitely many $U(N)$-homogeneous strata indexed by isotypic multiplicities, with stabilizers given by centralizers; the canonical two-level sector is organized by $\mathrm{Gr}_{2}(\mathbb{C}^{N})$ up to a $PSU(2)$ gauge. Equipping $U(N)$ with the Hilbert–Schmidt bi-invariant metric, each embedded subgroup is totally geodesic. Using two-level QR/Givens factorization together with an explicit generation of diagonal tori by two-level phase rotations, we prove phase-free universality $\langle\mathcal{G}^{SU}_{\mathrm{2lvl}}(n)\rangle=SU(N)$ and hence $\langle\mathcal{G}^{SU}_{\mathrm{elem}}(n)\rangle=SU(N)$. Full universality in $U(N)$ follows by adjoining the abelian diagonal/global $U(1)$ factors (equivalently, by passing to the $U(2)$ two-level dictionary). Finally, we record a modular finite-alphabet interface by lifting Solovay–Kitaev approximation in $SU(2)$ through two-level embeddings.
Keywords: quantum gates, Lie group embeddings, homogeneous manifolds, Grassmannian, Riemannian geometry, universality, Solovay–Kitaev theorem
@@ 1 Introduction
Main results as a single chain: landscape $\rightarrow$ two-level sector $\rightarrow$ universality and finite interface. The contributions of the paper are best read as a three-step chain.
(R1) The embedding landscape is computable. Using complete reducibility of unitary $SU(2)$-representations, we describe $\mathrm{Emb}(SU(2),U(N))$ as a disjoint union of $U(N)$-homogeneous manifolds indexed by irreducible multiplicity data, and we compute stabilizers explicitly as centralizers. This yields a finite orbit-type decomposition under conjugation by $U(N)$, together with closed-form dimension formulas. Conceptually, it turns the choice of an intrinsic “logical degree of freedom” inside $U(N)$ into discrete, computable data.
(R2) A canonical two-level sector provides an immediate model. Within the landscape, the two-level stratum is particularly transparent and serves as the main computational interface: choosing a logical qubit amounts to selecting a complex two-plane $W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})$. Elementary operations then act as $SU(2)$ on $W$ and as the identity on $W^{\perp}$. An internal $PSU(2)$ gauge captures the redundancy in choosing a frame on $W$, separating gauge-invariant support data from frame-dependent labels. This Grassmannian sector interfaces directly with two-level (Givens-type) synthesis and, later, with the finite-alphabet lifting pipeline.
(R3) Expressiveness and discretization follow already from the two-level sector. A viable intrinsic vocabulary must guarantee universality and admit compilation. Letting
$\mathcal{G}^{SU}_{\mathrm{2lvl}}(n):=\bigcup_{W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})}SU(N)[W],\qquad\mathcal{G}^{U}_{\mathrm{2lvl}}(n):=\bigcup_{W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})}U(N)[W],$
we prove constructive generation results
$\big\langle\mathcal{G}^{SU}_{\mathrm{2lvl}}(n)\big\rangle=SU(N),\qquad\big\langle\mathcal{G}^{U}_{\mathrm{2lvl}}(n)\big\rangle=U(N),$
and therefore, by inclusion $\mathcal{G}^{(\cdot)}_{\mathrm{2lvl}}(n)\subset\mathcal{G}^{(\cdot)}_{\mathrm{elem}}(n)$, the embedding-elementary dictionaries are universal as well. The proofs follow a classical QR/Givens elimination up to a diagonal special-unitary factor, together with an explicit generation of the diagonal torus by two-level one-parameter subgroups; the passage from $SU(N)$ to $U(N)$ is then obtained by explicit diagonal/global phase bookkeeping in the $U(2)$ extension, keeping the non-abelian synthesis cleanly separated from abelian phase management. Finally, we record a modular finite-alphabet interface: exact two-level factorizations reduce synthesis in $U(N)$ to repeated synthesis in $U(2)$ (or in $SU(2)$ after phase normalization); Solovay–Kitaev provides polylogarithmic approximation of $SU(2)$ elements by words over a fixed finite universal set; and these words lift canonically to $U(N)$ through two-level embeddings, with global operator-norm error control. Diagonal and global-phase factors may be carried explicitly or compiled separately, depending on the desired normal form.
Variational meaning (supporting layer). Equipping $U(N)$ with the Hilbert–Schmidt bi-invariant metric, we show that each embedded copy $\phi(G)\subset U(N)$ (with $G\in\{SU(2),U(2)\}$) is totally geodesic. Consequently, within a fixed embedded gate manifold, energy-minimizing implementations of a target element are exactly constant-speed one-parameter subgroups generated by minimal-norm logarithms in the embedded Lie algebra. In the two-level sector this identifies the intrinsic metric cost in $U(N)$ with the corresponding $2\times 2$ generator, providing a clean variational interpretation of elementary motions.
@@
Several ingredients used in this paper are classical when taken in isolation: the classification of finite-dimensional unitary $SU(2)$-representations and their commutants via isotypic decompositions and Schur-type arguments, and the compilation toolkit based on two-level (Givens/QR) factorizations and Solovay–Kitaev approximation in $SU(2)$. What is not standard is to treat these ingredients as components of a single intrinsic descriptor layer for gate primitives internal to $U(2^{n})$: an elementary operation is specified by the choice of a faithful embedded copy of $SU(2)$ (or $U(2)$ when phases are retained), together with its gauge structure and natural parameter spaces (most notably the Grassmannian two-level sector). This viewpoint yields, in one framework, (i) a computable stratification of admissible intrinsic two-level degrees of freedom, (ii) a canonical geometric model for “logical qubits” inside $U(N)$, and (iii) a modular compilation interface with uniform operator-norm control obtained by lifting finite $SU(2)$ approximation schemes through two-level embeddings.
@@ 3.5 Relation to the embedding dictionary and to two-level unitaries
$\mathcal{G}_{\mathrm{2lvl}}^{SU}(n)\;:=\;\bigcup_{W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})}SU(N)[W]\;\subset\;\mathcal{G}_{\mathrm{elem}}^{SU}(n),\qquad\mathcal{G}_{\mathrm{elem}}^{SU}(n)=\bigcup_{\phi\in\mathrm{Emb}(SU(2),U(N))}\phi(SU(2)).$
@@ 4.1 Diagonal tori from two-level phase rotations
Consequently, every $D\in T$ admits a factorization $D=e^{i\theta}D_{0}$ with $\theta\in\mathbb{R}$ and $D_{0}\in\langle\,SU(N)[W_{12}]\cup\cdots\cup SU(N)[W_{1N}]\,\rangle$.
@@ 4.3 Universality: phase-free and full
$\big\langle\mathcal{G}^{SU}_{\mathrm{elem}}(n)\big\rangle=SU(N),\qquad\big\langle\mathcal{G}^{U}_{\mathrm{elem}}(n)\big\rangle=U(N).$
@@ 5.5 Algorithmic summary and interpretation
6 The embedding landscape: $\mathrm{Emb}(G,U(N))$ as a union of homogeneous manifolds
@@ 6.4 Remark on $U(2)$ and determinant twists
7 Riemannian geometry and a variational principle for elementary gates
@@ 7.2 Bi-invariant geodesics and totally geodesic subgroups
For a bi-invariant metric, left and right translations are isometries. The Levi–Civita connection satisfies $\nabla_{X^{R}}Y^{R}=\tfrac{1}{2}[X,Y]^{R}$ for right-invariant fields, hence the geodesic equation is equivalent to constant body velocity. Thus $\gamma(t)=g_{0}\exp(tX)$. See [9, Ch. 4] or [14, § II.1]. ∎

### v2

# arXiv:2601.17936v2[quant-ph] 02 Mar 2026

@@ Abstract
In the standard circuit model, elementary gates are defined relative to a chosen tensor factorization and are therefore extrinsic to the ambient group $U(2^{n})$. Writing $N=2^{n}$, we introduce an intrinsic descriptor layer in $U(N)$ by declaring as primitive the motions inside faithful embedded copies of $SU(2)$ (phase-free), together with a phase-inclusive $U(2)$ variant. We describe the embedding landscape $\mathrm{Emb}(SU(2),U(N))$ as a finite union of $U(N)$-homogeneous strata indexed by isotypic multiplicities, with stabilizers given by centralizers, and we isolate a canonical two-level sector parameterized by $\mathrm{Gr}_{2}(\mathbb{C}^{N})$ up to a $PSU(2)$ gauge. Equipping $U(N)$ with the Hilbert–Schmidt bi-invariant metric, each embedded subgroup is totally geodesic, yielding a variational characterization of elementary motions via minimal-norm logarithms. On the constructive side, we prove phase-free universality in $SU(N)$ from two-level primitives using QR/Givens factorizations together with explicit diagonal generation, and we obtain full universality in $U(N)$ by explicit abelian phase bookkeeping (equivalently, via the $U(2)$ two-level dictionary). Finally, we formalize a modular finite-alphabet compilation interface: any approximation routine in $SU(2)$ (e.g. Solovay–Kitaev) can be lifted through two-level embeddings to yield $U(N)$-level synthesis with global operator-norm error control.
Keywords: quantum gate synthesis, quantum compiling, intrinsic gate descriptors, two-level unitaries (Givens/QR), $SU(2)$ embeddings in $U(2^{n})$, Grassmannian, homogeneous manifolds, Riemannian geometry, universality, Solovay–Kitaev
@@ 1 Introduction
Main results and technical narrative. We develop a single pipeline from an intrinsic definition of elementarity in $U(N)$ to a modular synthesis/compilation interface. At the structural level, we classify faithful $SU(2)$ embeddings in $U(N)$ up to conjugation, obtaining a finite orbit-type stratification indexed by isotypic multiplicities and explicit stabilizers (centralizers). At the computational level, we isolate the Grassmannian two-level sector (logical qubits as $W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})$ up to a $PSU(2)$ gauge), prove constructive universality from two-level primitives via QR/Givens factorizations plus explicit diagonal generation, and formalize a finite-alphabet interface by lifting $SU(2)$ approximation routines (e.g. Solovay–Kitaev) through two-level embeddings with global operator-norm error control.
Variational meaning. Equipping $U(N)$ with the Hilbert–Schmidt bi-invariant metric, we show that each embedded copy $\phi(G)\subset U(N)$ (with $G\in\{SU(2),U(2)\}$) is totally geodesic. Hence, within a fixed embedded gate manifold, energy-minimizing implementations of a target element are exactly constant-speed one-parameter subgroups generated by minimal-norm logarithms in the embedded Lie algebra. In the two-level sector this identifies the intrinsic metric cost in $U(N)$ with the corresponding $2\times 2$ generator, providing a direct variational interpretation of elementary motions.
@@
Several ingredients are classical in isolation: the representation theory of $SU(2)$, two-level (Givens/QR) synthesis, and finite-alphabet approximation in $SU(2)$ (e.g. Solovay–Kitaev). The contribution here is to organize these components into a single intrinsic, conjugation-invariant descriptor layer for elementary gates internal to $U(2^{n})$: primitives are motions inside faithful SU(2)SU(2)/$U(2)$ embeddings, equipped with a natural gauge structure and parameter spaces (notably the Grassmannian two-level sector), and this viewpoint yields a modular compilation interface obtained by lifting $SU(2)$ approximation routines through two-level embeddings with uniform operator-norm control.
Contributions (intrinsic formalism and compilation interface).
We introduce an intrinsic descriptor layer for elementary operations in $U(N)$ ($N=2^{n}$) by declaring as primitives the motions inside faithful embedded copies of $SU(2)$ (and its $U(2)$ variant), yielding phase-free and phase-inclusive dictionaries. We identify a canonical two-level sector organized by $\mathrm{Gr}_{2}(\mathbb{C}^{N})$ up to a $PSU(2)$ gauge, and we describe the embedding landscape as finitely many homogeneous strata indexed by isotypic multiplicities. We prove phase-free universality in $SU(N)$ from two-level primitives via QR/Givens factorization plus explicit diagonal generation; full universality in $U(N)$ follows by adjoining abelian diagonal/global factors (or by passing to the $U(2)$ two-level dictionary). We provide a modular finite-alphabet compilation interface: any $SU(2)$ approximation routine (e.g. Solovay–Kitaev) can be lifted through two-level embeddings with global operator-norm error control.
Scope and non-claims. We do not claim improved worst-case complexity bounds for gate synthesis, nor optimality of any particular compilation subroutine. Our aim is to supply an intrinsic, representation- and gauge-aware language for elementary gates in $U(2^{n})$ together with a clean, modular pipeline connecting $U(N)$-factorizations to $SU(2)$ finite-alphabet approximation.
Supporting layers (scope alignment). Sections 6 and 7 provide structural and interpretive backbones for the intrinsic descriptor layer, rather than additional compilation claims. The embedding-landscape analysis makes the admissible intrinsic two-level degrees of freedom computable (finite orbit types with explicit stabilizers/centralizers), while the variational section provides a clean baseline metric cost for elementary motions (total geodesy and minimal-logarithm characterizations) that can be replaced by more hardware-specific control models.
@@ 3.5 Relation to the embedding dictionary and to two-level unitaries
$\mathcal{G}^{SU}_{\mathrm{2lvl}}(n)$ $=\bigcup_{W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})}SU(N)[W],$
$\mathcal{G}^{U}_{\mathrm{2lvl}}(n)$ $=\bigcup_{W\in\mathrm{Gr}_{2}(\mathbb{C}^{N})}U(N)[W].$
@@ 4.1 Diagonal tori from two-level phase rotations
Consequently, every $D\in T$ admits a factorization $D=e^{i\theta}D_{0}$ with $\theta\in\mathbb{R}$ and $D_{0}\in\big\langle\,SU(N)[W_{12}]\allowbreak\cup\cdots\allowbreak\cup SU(N)[W_{1N}]\,\big\rangle$.
@@ 4.3 Universality: phase-free and full
$\big\langle\mathcal{G}^{SU}_{\mathrm{2lvl}}(n)\big\rangle$ $=SU(N),$
$\big\langle\mathcal{G}^{U}_{\mathrm{2lvl}}(n)\big\rangle$ $=U(N).$
@@ 5.5 Algorithmic summary and interpretation
5.6 Relation to fault-tolerant and hardware-constrained gate sets
The finite-alphabet interface recorded in this section is deliberately agnostic to the physical origin of the discrete gate set. In fault-tolerant settings, the available alphabet is typically restricted (e.g. Clifford+$T$-type libraries, lattice-surgery primitives, or other architecture-dependent native gates). Our framework separates (i) the intrinsic two-level lifting layer in $U(N)$ from (ii) the local approximation engine in $SU(2)$: any routine that approximates target elements of $SU(2)$ over a prescribed finite alphabet can be inserted into the pipeline, while the global error control after lifting through two-level embeddings remains unchanged.
Importantly, we do not claim optimal resource counts for any specific fault-tolerant metric (such as $T$-count). Rather, the point is that hardware- or FTQC-imposed alphabets can be incorporated without altering the intrinsic descriptor layer or the two-level universality backbone, thus providing a clean interface between $U(N)$-factorizations and architecture-specific synthesis.
6 Structural backbone: the embedding landscape $\mathrm{Emb}(G,U(N))$ as a union of homogeneous manifolds
@@ 6.4 Remark on $U(2)$ and determinant twists
7 Supporting layer: a variational (metric) cost for elementary gates
@@ 7.2 Bi-invariant geodesics and totally geodesic subgroups
For a bi-invariant metric, left and right translations are isometries. The Levi–Civita connection satisfies $\nabla_{X^{R}}Y^{R}=\tfrac{1}{2}[X,Y]^{R}$ for right-invariant fields. Hence the geodesic equation is equivalent to constant body velocity. Thus $\gamma(t)=g_{0}\exp(tX)$. See [9, Ch. 4] or [14, § II.1]. ∎